If y varies inversely with x and y=2/9 when x=1/3, what in the constant of variation?
Answer by nerdybill(7384)

If y varies inversely with x and y=2/9 when x=1/3, what in the constant of variation?
.
When you se "inverse variation" think:
y = k/x
.
plug in what was given to find the constant k:
2/9 = k/(1/3)
multiplying both sides by 3:
2/9 * 3= k
2/3= k
.
answer: 2/3